i need help with this word problem can anyone help me?
Soybean meal is 16% protein, COrnmeal is 8% portein. Howmany pounds of each should be mixed together in order to get a 320lb mixture that is 14% proetien?
Cornmeal__________pounds
Soybean____________pounds

Answer by ptaylor(2198)

Let x=amount of soybean meal needed
Then 320-x=amount of cornmeal needed
Now we know that the amount of pure protein in the soybean meal (0.16x) plus the amount of pure protein in the cornmeal (0.08(320-x)) has to equal the amount of pure protein in the final mixture (0.14*320). So our equation to solve:
0.16x+0.08(320-x)=0.14*320 simplify
0.16x+25.6-0.08x=44.8 subtract 25.6 from each side
0.16x+25.6-25.6-0.08x=44.8-25.6 collect like terms
0.08x=19.2 divide both sides by 0.08
x=240 lb----------------------amount of soybean meal needed
320-x=320-240=80 lb-----------------amount of cornmeal needed
CK
240*0.16+80*0.08=0.14*320
38.4+6.4=44.8
44.8=44.8

Hope this helps---ptaylor